# Parameter estimation in fluid flow models from undersampled frequency space data

### v1

@@ 3.2.2 Forward problem setup
$\bm{u}_{in}=\bm{u}_{profile}Uf(t)$
where $\bm{u}_{profile}$ is a flow profile provided by the solution of a Stokes problem in the domain, $U$ is a constant amplitude, and $f(t)$ defines the temporal profile as
@@
with $T=0.64$, $\beta=5$. The shape was determined empirically to match the shape of the flow rate in the phantom.
@@ 3.2.3 Inverse problem setup
We are reparameterizing such that $\bm{\theta}=\bm{\theta}^{0}\odot\bm{\nu}$ with $\bm{\theta}^{0}$ the initial guess for the parameters and the filter being applied to $\bm{\nu}$. This was done because the exponential reparameterization used above proved unstable in this case.
@@
$\bm{Y}^{0}-\bm{M}(t^{0})\odot\exp\left(i\bm{\phi}_{back}(t^{0}))\right)\odot\bm{S}$
@@
$\bm{\phi_{back}}=\angle\mathcal{F}^{-1}(\bm{Y}_{back})$ (17)
where $\bm{Y}_{back}$ is a zero-filled measurement acquired with $R=2$ and no encoding gradient, i.e. capturing only the background magnetization and no fluid velocity.
@@
$\bm{M}=|\mathcal{F}^{-1}(\bm{Y})|$ (18)
for each velocity direction, where $\bm{Y}$ is a zero-filled measurement acquired with $R=2$ in that velocity direction. This however assumes that a highly sampled measurement is available. Therefore the other option is to use
$\bm{M}=|\mathcal{F}^{-1}(\bm{Y}_{back})|$ (19)
@@
The original data with an acceleration factor of 2 were acquired with an incoherent pseudo-spiral sampling in the $k_{y}-k_{z}$ direction with a fully sampled $k_{x}$-direction. We further undersample this by applying masks with a Gaussian or spiral sampling pattern onto the sampling mask for $R=2$ to achieve higher acceleration factors. In contrast to the masks for the synthetic data, this results in different masks for each time step and velocity direction. Examples of the resulting masks are shown in Figure 7.
The measurements are assumed to be placed at every $0.053s$. For the inverse problem, the first two measurements (at $0.053s$ and $0.106s$) are omitted as the presence of negative velocity values in these measurements could interfere with the Kalman filter, since this could lead to negative values in the particles which would be unphyscial for the forward simulation.
@@
(b) Spiral, $R=8$
(c) Spiral, $R=16$
(d) Spiral, $R=32$
(e) Spiral, $R=64$
(f) Spiral, $R=128$
(g) Gaussian, $R=8$
(h) Gaussian, $R=16$
(i) Gaussian, $R=32$
(j) Gaussian, $R=64$
(k) Gaussian, $R=128$
@@ 4.1 Synthetic data
In order to accurately reflect the different impact of the parameters on the flow, we compute the error in terms of the flow reconstructed from the estimated parameters. Using the estimated parameters, we solve the forward problem 14 again with a time step of $dt=0.001$. The error is calculated as
$e=\frac{||\bm{u}_{ref}-\bm{u}_{recon}||_{2}}{||\bm{u}_{ref}||_{2}}$
where $\bm{u}_{recon}$ and $\bm{u}_{ref}$ are vectors consisting of the velocity values for all components at each point in the geometry at each point in time ($dt_{meas}=15ms$) stacked together, using the estimated parameter values for $\bm{u}_{recon}$ and the true parameter values for $\bm{u}_{ref}$.
We compare the results of our method (directly estimating from frequency-space measurements) to velocity measurements reconstructed with compressed sensing using BART with a $l1$-regularization in time. The regularization parameter was determined empirically by visual evaluation for each combination of acceleration factor and mask and is listed in Table 3. For the estimation, we are using the method presented in [9] to take advantage of knowing the $R=2$ magnitude.
@@
Figure 8: Error values for different acceleration factors and different masks. Dotted lines are the error values of the CS-reconstructed velocity values themselves, dashed lines for the flow reconstructed from the inverse problem from BART velocities, solid lines are from the inverse problem in frequency space. Low initial guess on the left, high initial guess on the right. The bars indicate the standard deviation of the error.
The error values for all different acceleration factors for the low and high initial guess are depicted in Figure 8, for our method as well as using BART, as well as the error values of the reconstructed velocity itself (without first solving the inverse problems). In all cases, using the inverse problem to retrieve parameters and then reconstruct the flow achieves better results than the velocity measurements themselves. For the k-space cost function, the errors of both masks are very close to the error with fully sampled data for $R=8$, and increase for $R=16$ and $R=32$. The gaussian mask achieves lower error values for all three acceleration factors.
@@
By comparing the estimated parameter values of each parameter individually, as seen in Figure 9, it is apparent that the success of the estimation differs depending on the parameter. The inflow $U$ is estimated relatively accurately by all masks, with only a small decrease with increasing $R$. In comparison, the distal resistance of the first Windkessel outlet ($R_{d,1}$) is underestimated considerably and increasingly by the spiral mask, while the gaussian mask estimates it well for $R=8$ and $R=16$, but overestimates it severely for $R=32$. $R_{d,2}$ and $R_{d,3}$ show the same pattern, with the spiral mask underestimating the values compared to the gaussian mask, though more severely for $R_{d,2}$ than $R_{d,3}$. For $R=32$, the spiral mask slightly overestimates the value of $R_{d,3}$.
@@
We also consider the case where measurements are only available for one of the velocity components, in this case the $z$-component, which is equivalent to the foot-head direction in this setup. Figure 10 shows the error values for this case. The error values show the same patterns as when using all velocity components, and are only slightly higher. Here there is a lesser difference in the error values between the spiral and gaussian mask for $R=8$ and $R=16$, and the spiral mask shows a flatter error curve, whereas the error for the gaussian mask increases with $R=32$. Again, using the frequency measurements achieves lower error values than BART measurements for all subsampled data.
@@
Robustness to the choice of $venc$
As described previously in Remark 2, a low $venc$ leads to a high signal-to-noise ratio, but can lead to aliasing artifacts if the actual maximal velocity exceeds the $venc$. The cost function used in [9] remedies this by distinguishing the actual velocity from the wrapped ones when the physical parameters affect the velocity on several voxels simultaneously. In the present work, the cost function corresponds to a similar cost function as in [9] – but with an additional Fourier transform – therefore also including the aliasing compensation.
To investigate the robustness of our method to choosing $venc$s lower than the maximal velocity, thus utilizing the higher sensitivity to the velocity in the signal, we compare three additional $venc$ values corresponding to $80\%$, $30\%$, and $10\%$ of the maximum velocity. The results are shown in Figure 12.
@@
Figure 12: Error values for different values of the $venc$. Spiral mask on the left, Gaussian mask on the right. The value for $R32$ for the spiral mask for the lowest venc is excluded as the inverse problem failed to converge. For the BART reconstructions, $venc=10\%$ was not considered.
The results improve considerably with a decreasing $venc$ for 80% and 30%, especially for higher acceleration factors. However, the smallest $venc$ of 10% leads to increased error values and standard deviations, with the inverse problem even failing to provide results for the spiral mask for $R=32$. The results for both masks perform similarly with the decreasing $venc$, indicating that this response is inherent to the inverse problem formulation – as shown in [9] – rather than a result of the choice of mask.
In comparison, with the BART reconstructions, the error remains similar for $venc=80\%V_{max}$ and then increases significantly for $venc=30\%V_{max}$. This behaviour can be explained from the fact that the reconstructed velocities were already unwrapped prior to the parameter estimation as described in [9].
@@
We also provide examples of the parameter curves of ROUKF in Figure 14, which show the development of the estimated parameter over the length of the simulation. The shaded area indicates the variance of the parameter. This parameter variance, for each of the masks using our new method, is higher at the end of the simulation for higher $R$, since less data is available. This effect is not observed for the BART-reconstructed velocities, where the variance decreases similarly for all acceleration factors. Here the variance is also much smaller than for the frequency-space based method, indicating higher confidence in the estimated parameter. This higher confidence is based on the fact that due to the reconstruction, the measurements contain more data than in frequency space (as they now have as many voxels as the image space, as compared to only the actually measured voxels). Since these data however originate from the same subsampled measurements, they do not actually contain more information, which leads to this misleading quantification of the uncertainty in the inverse problem.
Furthermore, it can be seen that for low $R$, the curves qualitatively follow the shape of the curve for the fully sampled data, but with higher sampling rates, the choice of the mask not only quantitatively but also qualitatively impacts the evolution of the parameter as it is estimated. The curves from our method also differ considerably from those using BART, which develop erratic spikes in either direction as the acceleration factor increases. In comparison to that, the curves for each mask remain more similar to each other with increasing $R$.
@@ 4.2 Phantom data
The results of the estimation of the noise per mask and acceleration factor are shown in Figure 15. While the estimation of the noise with the spiral mask remains close to constant for the acceleration factors, the estimations with the gaussian mask show major variations for some of the coils. This is likely due to the spatial variation of the signal strength for some of the coils and the fact that the gaussian mask does not uniformly cover the space. Nonetheless this approach provides a good approximation of the standard deviation of the noise even for high acceleration factors.
@@
We are using the same error metric as described in Section 4.1, with $\bm{u}_{ref}$ the solution of a forward problem with the parameters estimated from $R=2$.
Figure 16: Error values for different acceleration factors for both error metrics. The ”mag0” labelled results use the magnitude from the acquisition without a velocity encoding gradient.
The error values can be seen in Figure 16. A considerable difference between the behaviour of the two masks is apparent. Both achieve similar results for an acceleration factor of $R=16$, but the error increases higher for the spiral mask than for the Gaussian mask. The Gaussian mask also shows a decrease in error from $R=64$ to $R=128$, which may be a result of differences in the noise estimation, or otherwise an outlier. Both choices for which magnitude to use in the observation show similar results, implying robustness to using a magnitude from a different velocity direction.
In Figure 17, we show the flow rates of the reconstructed flows at the inlet and each of the outlets. It can be seen that both masks tend to overestimate the inflow with increasing acceleration factor $R$. For the gaussian mask, the flow split between the two outlets remains fairly consistent as $R$ increases, whereas the spiral mask no longer accurately depicts the flow split from $R=32$ upwards. This can also be observed by looking at the estimated parameter values in Figure 18. Both masks show a very consistent estimation of the inflow amplitude $U$, with differences only for $R=128$, but the estimation of the resistance boundary condition $R_{p,2}$ shows larger changes for different acceleration factors. The spiral mask especially overestimates $R_{p,2}$ for higher acceleration factors, leading to the lack of distinction of the outlets in the flow rates.
@@ 5 Discussion
The numerical results are consistent between the simulated dataset and the phantom dataset. In both cases, our approach robustly estimates boundary conditions of the flow even with high undersampling rates and low $venc$ values and outperforms the conventional approach of reconstructing velocity measurements from undersampled data through Compressed Sensing.
The results also show significant differences between the two studied sampling patterns, with qualitatively and quantitatively different outcomes of the inverse problem. The sampling patterns also appear to differ in their estimation of different parameters, indicating that the choice of sampling pattern may depend on which parameter should be estimated more accurately.
A limitation of the method is that it requires an accurate estimate of the magnitude and the background phase of the magnetization. However, since the magnitude can be scanned without motion encoding gradients, a highly sampled image of the magnitude is quick to acquire. We have also shown that it was possible – at least in the phantom dataset – to acquire the background image highly sampled for approximations of the magnitude and background phase, while undersampling the remaining three directions. As we have shown some robustness of the method with regards to flaws in the magnitude, one could also acquire the magnitude only at diastole, and use this for all time steps to cut down on acquisition time, or correct it using the model itself. Additionally, the noise present in the reconstructed magnitude and background phase is currently not considered by the inverse problem. This will be subject of future research.
Another limitation of this paper is that real patient data is not included, which would pose additional complexities such as the movement of the vessels due to the breathing of the patient.

### v2

@@ 3.2.2 Forward problem setup
$\boldsymbol{u}_{in}=\boldsymbol{u}_{profile}Uf(t)$
where $\boldsymbol{u}_{profile}$ is a flow profile provided by the solution of a Stokes problem in the domain, $U$ is a constant amplitude, and $f(t)$ defines the temporal profile as
@@
with $T=0.64$, $\beta=5$. The temporal profile was determined empirically to match the shape of the flow rate in the phantom.
@@ 3.2.3 Inverse problem setup
We are reparameterizing such that $\boldsymbol{\theta}=\boldsymbol{\theta}^{0}\odot\boldsymbol{\nu}$ with $\boldsymbol{\theta}^{0}$ the initial guess for the parameters and the filter being applied to $\boldsymbol{\nu}$, with the initial value of $\boldsymbol{\nu}=\boldsymbol{1}$. This was done because the exponential reparameterization used before proved unstable in this case.
@@
$\boldsymbol{Y}^{0}-\boldsymbol{M}(t^{0})\odot\exp\left(i\boldsymbol{\phi}_{back}(t^{0}))\right)\odot\boldsymbol{S}$
@@
$\boldsymbol{\phi_{back}}=\angle\mathcal{F}^{-1}(\boldsymbol{Y}_{back})$ (17)
where $\boldsymbol{Y}_{back}$ is a zero-filled measurement acquired with $R=2$ and no encoding gradient, i.e. capturing only the background magnetization and no fluid velocity.
@@
$\boldsymbol{M}=|\mathcal{F}^{-1}(\boldsymbol{Y})|$ (18)
for each velocity direction, where $\boldsymbol{Y}$ is a zero-filled measurement acquired with $R=2$ in that velocity direction. This however assumes that a highly sampled measurement is available. Therefore the other option is to use
$\boldsymbol{M}=|\mathcal{F}^{-1}(\boldsymbol{Y}_{back})|$ (19)
@@
The original data with an acceleration factor of 2 were acquired with an incoherent pseudo-spiral sampling in the $k_{y}-k_{z}$ direction with a fully sampled $k_{x}$-direction. We further undersample this by applying masks with a Gaussian or spiral sampling pattern onto the sampling mask for $R=2$ to achieve higher acceleration factors $R=16,32,64,128$. In contrast to the masks for the synthetic data, this results in different masks for each time step and velocity direction. Examples of the resulting masks are shown in Figure 7.
The measurements are assumed to be placed at every $0.053s(\approx 1s/19)$. For the inverse problem, the first two measurements (at $0.053s$ and $0.106s$) are omitted as the presence of negative velocity values in these measurements could interfere with the Kalman filter, since this could lead to negative values in the particles which would be unphyscial for the forward simulation.
@@
(b) Spiral, $R=16$
(c) Spiral, $R=32$
(d) Spiral, $R=64$
(e) Spiral, $R=128$
(f) Gaussian, $R=16$
(g) Gaussian, $R=32$
(h) Gaussian, $R=64$
(i) Gaussian, $R=128$
@@ 4.1 Synthetic data
Computational effort. The parameter estimation with either velocity data or frequency data requires the same amount of CPU time (around 13 minutes using an AMD 7763 CPU), though it can be run in parallel to reduce the walltime. Using velocity measurements uses slightly less memory however, as no complex-value data have to be saved (888.6MB compared to 959MB for the frequency-space data). Performing the reconstruction of the velocity data with BART requires an additional 1:30min of CPU time and 660MB of memory.
We compare the results of our method (directly estimating from frequency-space measurements) to estimating parameters from velocity measurements reconstructed with compressed sensing using BART with a $l1$-regularization in time. The regularization parameter was determined empirically by visual evaluation for each combination of acceleration factor and mask and is listed in Table 3. For the estimation, we are using the method presented in [garay2022] to take advantage of knowing the $R=2$ magnitude.
Comparing the reconstruted flows. In order to accurately reflect the different impact of the parameters on the flow, we compute the error in terms of the flow computed from the estimated parameters. Using the estimated parameters, we solve the forward problem 14 again with a time step of $dt=0.001$. The error is calculated as
[…]
where $\boldsymbol{u}_{recon}$ and $\boldsymbol{u}_{ref}$ are vectors consisting of the velocity values for all components at each point in the geometry at each point in time stacked together, using the estimated parameter values for $\boldsymbol{u}_{recon}$ and the true parameter values for $\boldsymbol{u}_{ref}$.
@@
Figure 8: Error values for different acceleration factors and different masks. Dotted lines are the error values of the flow reconstructed from the parameters estimated by the inverse problem using zero-filled measurements, dashed lines for the flow reconstructed from the inverse problem from BART velocities, solid lines are from the inverse problem in frequency space. Low initial guess on the left, high initial guess on the right. The bars indicate the standard deviation of the error.
The error values for all different acceleration factors for the low and high initial guess are depicted in Figure 8, for our method as well as for estimating the parameters using velocities reconstructed with BART, as well as the error values of the velocity reconstructed with BART itself (without first estimating the parameters). In the latter case, $u_{ref}$ was interpolated into the coarser mesh to match the resolution of the reconstructed velocities. In all cases, estimating the parameters and then reconstructing the flow achieves better results than using zero-filled velocity measurements. For the k-space cost function, the errors of both masks are very close to the error with fully sampled data for $R=8$, and increase for $R=16$ and $R=32$. The Gaussian mask achieves lower error values for all three acceleration factors.
@@
Comparing the parameter values. By comparing the estimated parameter values of each parameter individually, as seen in Figure 9, it is apparent that the success of the estimation differs depending on the parameter. The inflow $U$ is estimated relatively accurately by all masks, with only a small decrease with increasing $R$. In comparison, the distal resistance of the first Windkessel outlet ($R_{d,1}$) is underestimated considerably and increasingly by the spiral mask, while the gaussian mask estimates it well for $R=8$ and $R=16$, but overestimates it severely for $R=32$. $R_{d,2}$ and $R_{d,3}$ show the same pattern, with the spiral mask underestimating the values compared to the Gaussian mask, though more severely for $R_{d,2}$ than $R_{d,3}$. For $R=32$, the spiral mask slightly overestimates the value of $R_{d,3}$.
@@
Using only a single velocity component. We also consider the case where measurements are only available for one of the velocity components, in this case the $z$-component, which is equivalent to the foot-head direction in this setup. Figure 10 shows the error values for this case. The error values show the same patterns as when using all velocity components, and are only slightly higher. Here there is a lesser difference in the error values between the spiral and Gaussian mask for $R=8$ and $R=16$, and the spiral mask shows a flatter error curve, whereas the error for the Gaussian mask increases with $R=32$. Again, using the frequency measurements achieves lower error values than BART measurements for all subsampled data.
@@
Robustness to the choice of vencvenc
As described previously in Remark 2, a low vencvenc leads to a high signal-to-noise ratio, but can lead to aliasing artifacts if the actual maximal velocity exceeds the vencvenc. The cost function used in [garay2022] remedies this by distinguishing the actual velocity from the wrapped ones when the physical parameters affect the velocity on several voxels simultaneously. In the present work, the cost function corresponds to a similar cost function as in [garay2022] – but with an additional Fourier transform – therefore also including the aliasing compensation.
To investigate the robustness of our method to choosing vencvencs lower than the maximal velocity, thus utilizing the higher sensitivity to the velocity in the signal, we compare three additional vencvenc values corresponding to $80\%$, $30\%$, and $10\%$ of the maximum velocity. The results are shown in Figure 12.
@@
Figure 12: Error values for different values of the vencvenc. Spiral mask on the left, Gaussian mask on the right. The value for $R32$ for the spiral mask for the lowest venc is excluded as the inverse problem failed to converge. For the BART reconstructions, $venc=10\%$ was not considered.
The results improve considerably with a decreasing vencvenc for 80% and 30%, especially for higher acceleration factors. However, the smallest vencvenc of 10% leads to increased error values and standard deviations, with the Kalman filter even failing to provide results for the spiral mask for $R=32$. The results for both masks perform similarly with the decreasing vencvenc, indicating that this response is inherent to the inverse problem formulation – as shown in [garay2022] – rather than a result of the choice of mask.
In comparison, with the BART reconstructions, the error remains similar for $venc=80\%V_{max}$ and then increases significantly for $venc=30\%V_{max}$. This behaviour can be explained from the fact that the reconstructed velocities were already unwrapped prior to the parameter estimation as described in [garay2022].
@@
Parameter curves of the Kalman filter. We also provide examples of the parameter curves of ROUKF in Figure 14, which show the development of the estimated parameter over the length of the simulation. The shaded area indicates the variance of the parameter. This parameter variance, for each of the masks using our new method, is higher at the end of the simulation for higher $R$, since less data is available. This effect is not observed for the BART-reconstructed velocities, where the variance decreases similarly for all acceleration factors. Here the variance is also much smaller than for the frequency-space based method, indicating higher confidence in the estimated parameter. This higher confidence is based on the fact that due to the reconstruction, the measurements contain more data than in frequency space (as they now have as many voxels as the image space, as compared to only the actually measured voxels). Since these data however originate from the same subsampled measurements, they do not actually contain more information, which leads to this misleading quantification of the uncertainty in the inverse problem.
Furthermore, it can be seen that for low $R$, the curves qualitatively follow the shape of the curve for the fully sampled data, but with higher sampling rates, the choice of the mask not only quantitatively but also qualitatively impacts the evolution of the parameter estimation. The curves from our method also differ considerably from those using BART, which develop erratic spikes in either direction as the acceleration factor increases. In comparison to that, the curves for each mask remain more similar to each other with increasing $R$.
@@
In summary, for the synthetic data,
  • using the inverse problem in frequency space significantly reduces the error
  • there are notable differences in the qualitative and quantitative estimation of the parameters between the different masks
  • the method is robust to different values of the vencvenc and using estimated magnitudes and background phases of the data.
@@ 4.2 Phantom data
First, we consider the quality of the estimation of the standard deviation of the noise in the case of the phantom data. The results per mask and acceleration factor are shown in Figure 15. While the estimation of the noise with the spiral mask remains close to constant for the acceleration factors, the estimations with the Gaussian mask show major variations for some of the coils. This is likely due to the spatial variation of the signal strength for some of the coils and the fact that the Gaussian mask does not uniformly cover the space. Nonetheless this approach provides a good approximation of the standard deviation of the noise even for high acceleration factors.
@@
For comparing the reconstructed flows, we are using the same error metric 20 as in Section 4.1, with $\boldsymbol{u}_{ref}$ the solution of a forward problem with the parameters estimated from $R=2$.
Figure 16: Error values for different acceleration factors for both masks. The ”mag0” labelled results use the magnitude from the acquisition without a velocity encoding gradient.
The error values can be seen in Figure 16. A considerable difference between the behaviour of the two masks is apparent. Both achieve similar results for an acceleration factor of $R=16$, but the error increases higher for the spiral mask than for the Gaussian mask. The Gaussian mask also shows a decrease in error from $R=64$ to $R=128$, which may be a result of differences in the noise estimation, or otherwise an outlier. Both choices for the magnitude in the observation show similar results, implying robustness to using a magnitude from a different velocity direction.
In Figure 17, we show the flow rates of the reconstructed flows at the inlet and each of the outlets. It can be seen that both masks tend to overestimate the inflow with increasing acceleration factor $R$. For the Gaussian mask, the flow split between the two outlets remains fairly consistent as $R$ increases, whereas the spiral mask no longer accurately depicts the flow split from $R=32$ upwards. This can also be observed by looking at the estimated parameter values in Figure 18. Both masks show a very consistent estimation of the inflow amplitude $U$, with differences only for $R=128$, but the estimation of the resistance boundary condition $R_{p,2}$ shows larger changes for different acceleration factors. The spiral mask especially overestimates $R_{p,2}$ for higher acceleration factors, leading to a lack of distinction of the outlets in the flow rates.
@@
In summary, for the phantom data,
  • the method successfully estimates the parameters with time-dependent and estimated magnitudes, background phases, and noise level
  • using the gaussian mask, the estimation is reasonably accurate even for very high acceleration factors (R=128)
  • the results are quantitatively and qualitatively different for the different masks, with the Spiral mask failing to distinguish between the two outlets for higher acceleration factors.
@@ 5 Discussion
The numerical results are consistent between the synthetic dataset and the phantom dataset. In both cases, our approach robustly estimates boundary conditions of the flow even with high undersampling rates and low vencvenc values and outperforms the conventional approach of reconstructing velocity measurements from undersampled data through Compressed Sensing.
The results also show significant differences between the two studied sampling patterns, with qualitatively and quantitatively different outcomes of the inverse problem. The sampling patterns appear to differ in their estimation of different parameters, indicating that the choice of sampling pattern may depend on which parameter is of higher interest.
In related research, in [muller_reduced-order_2018], Fourier-transformed time-varying measurements (for only a single spatial point) to estimate parameters of a one-dimensional fluid flow model, but this approach (neither the model nor the inverse problem) does not utilize the potential of using velocity data of the entire space in the frequency domain. Additionally, [zhao_direct_2016] used an Unscented Kalman Filter with undersampled frequency-space data to retrieve T2 mappings for MRI brain scans. However, they are relying on an assumption that the phase is both known and constant, hence only utilizing the magnitude of the k-space images, as well as estimating the state rather than a number of parameters. Additionally, they only experimented with relatively small acceleration factors (up to a factor of 8).
A limitation of the method is that it requires an accurate estimate of the magnitude and the background phase of the magnetization. However, since the magnitude can be scanned without motion encoding gradients, a highly sampled image of the magnitude is quick to acquire. We have also shown that it was possible – at least in the phantom dataset – to acquire the background image highly sampled for approximations of the magnitude and background phase, while undersampling the remaining three directions. As we have shown some robustness of the method with regards to flaws in the magnitude, one could also acquire the magnitude only at diastole, and use this for all time steps to cut down on acquisition time, or correct it using the model itself. Additionally, the noise present in the reconstructed magnitude and background phase is currently not considered by the Kalman filter. This will be subject of future research.
Another limitation of this paper is that subject data is not included, which would pose additional complexities such as the respiratory motion of the vessels due to the breathing of the patient and the potential large size of patient-specific meshes. We plan to address this aspect in future research.